Blog

Imaging, Single and Dual Aperture

The evenings of Feb. 12 and 13 were spent executing several different observing programs with LMIRCam (and some supporting NOMIC data). Winds were high and seeing mediocre on Feb. 12. It was much more reasonable on Feb. 13. We are now preparing for our last night of Nulling. We have a new algorithm to try and plenty of sources lined up to observe for the night.

4 comments

Second February Nulling Night: First Nulling Disk Detection

We saw our first dust disk last night with the LBTI nulling interferometer! First, the night details:

We carried out further nulling engineering last night, along with some test science to determine the state of the system. We began the night by optimizing our group delay tracker (also called the contrast gradient approach). This has allowed us to stabilize the phase, but has higher RMS variation than we need to hit our specs. We tried adding an integral term to this, and saw no improvement. Based on simulatons from Amali and Bertrand, we also tried merging the group delay and phase delay, to fine tune the performance. Also no improvement there. Its good to have ruled these out as the cause of the large RMS, but was disappointing.

We made more progress on scripting last night. Denis now has an automated routine for setting the group delay setpoint and routine. Dual position nodding was also further refined.

We took test data on fainter stars last night (1.7 Jy) with good results. This will be important for carrying out observations on the full HOSTS sample.

Finally, we looked at a star with a bright dust disk to test our performance. Denis carried out on-the-fly data reduction which confirmed we were seeing the disk. It made for a good end of the night.

Denis shows off the first LBTI disk detection
Denis shows off the disk detection.

Bill and Rafael run the AO systems
Bill and Rafael run the AO systems.

Our first detection. (rough reduction only).





10 comments

First February Nulling Night

Last night was our first night testing out the nulling and phasing system since December. Our first three nights, used for general commissioning and science, were rocky and plagued by poor seeing. We had several failures which has made working with the system more challenging. Our right fine alignment mirror stopped working, presumably due to a problem with cooling it down. Similarly, our right pupil alignment motor was found to have failed. These problems required us to warm and open the beamcombiner in order to fix the problems. On the AO side we are seeing an unusually high occurence of loop failures, which were affecting our productivity.

Sunday night started out similarly. However, by 11 pm or so the seeing had calmed down and the AO systems stopped misbehaving. In brief, we achieved several milestones:
  • Automated measurement of the null depth for use in scripts.
  • First testing of measuring the K band to N band setpoint automatically via a script
  • Measured and reduced the intensity mismatch to reduce the effect by 5x (~1% to 0.2%)
  • First testing of automated nodding via a script to allow nulling measurements in both beams (doubles our efficiency).
  • Acquired lengthy dataset to measure setpoint variations over time and elevation.
We are now on the sky for a second night and have more commissioning planned.

6 comments

February '14 Observing Run Weather Predictions

Observing Dates: February 7 - February 14.

0 comments

December 2013 Run Completed

Last night was the final observing night of the December 2013 observing block. In a nutshell, here's how it went. More details soon.



2 comments

December 2013 Run: First Stabilized Fringes

We have been doing a lot of testing. The last time I wrote we had confirmed that the phase sensing was working. However, control of it was not.

Last night we were able to stabilize fringes for the first time. The key to it was tracking the group delay. For the implementation we have this corresponds to a left-right gradient in the visibility. The movie below shows the loop grabbing the fringes and centering the fringe packet.



This allowed us to reduce the variations in the pathlength and set the 10 micron interference to be approximately at null (destructive interference). It also made people in the control room very happy.

We are about to go back on sky for our last night, so the details will need to wait for a later post. We are excited to start using our new 23 m telescope.

0 comments

December 2013 Run: Christmas day

Christmas day 2013, 20:00 pm -- the day and night crews having Christmas dinner at the LBT. Thanks to the participation of everyone, we had a delicious (and copious) meal including ham, asparagus, mashed potatoes, mac & cheese, "french" oignon beans, and various desserts.

Meanwhile in the control room, science observations are being carried out for LEECH. The telescope is nodding back and forth while LMIRcam is taking data. One might almost think that this is going on without any supervision....


But of course not. Andy has ingeniously put his laptop in front of the LMIRcam quick-look display and called Denis on skype.


Vanessa and Andy can now keep an eye on the observations from the dining room while we all share a good meal together on this special day.


We wish you a Merry Christmas from the LBT.


0 comments

December 2013 Run: Photo Blog

Hi LBTI Blog followers!

My name is Kim, and I'm a 3rd year graduate student who has been helping with the LBTI run over the past few days (along with post-doc Rob De Rosa and soon fellow grad student Abhi Rajan -- we all work in Jenny Patience's group at ASU). I've taken some photos while I've been here, and Phil and Vanessa suggested I write a guest post and add these to the blog. Hope you enjoy!

After this run started with a wintry blast, the skies finally cleared yesterday and began to melt the snow -- a good chance to snap a panorama from the balcony.

Cold and clear!

The snow survived long enough during the day for a snowy nighttime panorama too, with the city of Willcox visible on the left, along with the light domes of Tucson (middle) and Phoenix (rightmost). The starry skies belie an extraordinarily turbulent atmosphere -- at some points, we had seeing near 4 arcseconds!

4" seeing - not great for observing, but okay for panoramas. The Vatican Advanced Technology Telescope and the Sub-Millimeter Telescope can be seen on the right.
Much of the snow melted this morning and afternoon. By sunset, the forecast boded well for clear skies and improved seeing.

No inversion layer today. Clear as far as the eye can see.

With the moon approaching 3rd quarter, we've had the nice privilege of dark time for our first half nights. Below, the Milky Way provides a great backdrop for the open LBT dome. (Not pictured: IR starlight from a LEECH target makes its way into LMIRCam.)

Not sure what's happening with that little red light, sorry Andy!
Viewed from another angle, the Andromeda Galaxy (seen at the top left) and the Pleiades (top right) grace the LBT dome, while more stars within the galactic plane peek through the ventilation doors.

M31 as a small fuzzy - LBT - M45 (Pleiades)

(Also, a huge thanks to Chick Woodward for his kind donation of handwarmers, without which these nighttime shots would've been a lot less pleasant!)

Toward the northwest, the Milky Way sets over the treeline and surrounding forest.

A bit of light glow from Phoenix.
On the opposite horizon, the eastern sky rose bearing Jupiter (reflecting brightly on the left), Orion, and a bit more of the galactic plane in between. We'd soon be pointing the telescope in that direction to get a glimpse of the Galilean moons. The excellent dark sky at Mount Graham provides a nice view of M42, the Orion Nebula, to boot.

Spectacular dark skies - Jupiter looks brilliant.
Meanwhile, indoors we were getting some great results with excellent seeing (~0".5) and (mostly) nicely-behaved AO systems. The black plot on the middle monitor shows AO-corrected rms wavefront error of less than 1 nanometer on the highest orders, with overall correction (white curve) at a factor of 5-10 improvement over the input wavefronts (red curve) -- the first time we've reached these levels of correction. Vanessa approves!

Stamp of approval!

On the LMIRCam side, Andy and Phil pored over interferometric imaging at M' band of Jupiter's extremely volcanically-active moon, Io. Warm bright spots corresponding to volcanic activity could be seen on the moon's resolved disk.

Imaging Io with LMIRCAM.
The telescope phased well over the separation, providing great fringes of the bright spots!

Zooming in on the volcano spot.
Meanwhile on NOMIC, Denis was also spotting fringes of the Io volcano spots, this time at N-prime and 8.7 µm.

Volcano spots look great at longer wavelengths too!
Back outside, the cold clear night continued with reasonably good seeing, and a nice chance to get some parting star trails. A great end to a fantastic opportunity to work at the LBT and see the awesome progress being made with LBTI -- we've learned so much this run!

~20 minutes of exposure provides some small trails.








2 comments

December 2013 Run: Day 4

We finally got sky time tonight! We were able to open at sunset and begin observing soon after. The cold weather and difference in temperature causes some early night problems with AO stability (we were seeing large tip-tilt variations). The left AO system seems to operate fine on sky.

We scanned for fringes and had a system ready for phase testing by 9 pm. Seeing was okay early on, but degraded throughout the night.

We were again able to show regular control of the tip-tilt loop, but no significant improvement in stability with the phase loop. We did show, however, that for the first time we could adjust the phase setpoint of the system and have the corrector react as expected.

We identified several areas to test further:
  • Phase unwrapping was done only for a single wavelength jump. Multiple lambda jumps are now implemented.
  • Variations in tip-tilt may have been affecting phase stability, a work around was implemented.
  • poor SNR data may have been causing wavelength jumps. Revised software was implemented to remove these.
Attempts to test these variations later in the night were hampered by 1.5-4" seeing. We plan to continue testing of the phase loop after Christmas.



0 comments

December 2013 Run: Day 3

Another snow day on Saturday. We spent the time playing around with the test source. In particular, we explored vibration contributions. What we found:
  • Our pulse tube cooler does not contribute to vibrations seen with the artificial source
  • LUCI's coolers do not contribute to vibration.
  • Walking around on the back of the platform creates lots of vibrations.
  • Taping down the fold mirrors to dampen their motion doesn't reduce vibrations.
  • The cap. sensors on the FPC were creating very slow (3-10 Hz) responsiveness. We turned them off
Denis, Elwood, and Andy also refined the tip-tilt loop to use subpixel centroiding. This improved the performance. They are exploring an alternative phase calculation (parametric fitting, rather than angle of the Fourier Transform).

Sunday afternoon Vanessa and Phil realigned the SX wavefront sensor to allow removal of shims on the left dichroic. This should fix the vacuum leak we had earlier.

This afternoon it's been sunny but cold, so we are waiting to hear whether we can open tonight.



0 comments

December 2013 Run: Day 2

On friday the snow started. It has not been a very heavy snowfall, but has been pretty continuous. Day 2 was snowed out and we expect Day 3 to be lost as well.



In the meantime, Vanessa aligned the retroreflector on the left side and tested the AO system. She gave it all sorts of abuse (large tilts, simulated turbulence, calling it names) but the system showed no sign of the failure we saw in November. This is good news,and we are confident the left AO system is ready to go.

The AO GUI showing a big tilt on the mirror, right before the safety system takes over. This worked as expected.

Denis, Elwood, Andy, and Phil worked on realignment of the beamcombiner in the afternoon and evening. We slightly adjusted the ellipse on the left side to better align the beamcombiner to the telescope. A loose optic was identified and fixed on the right side. The beamcombiner artificial source was then able to be aligned.

In the evening we scanned (and scanned) for fringes. The optics have to have equal pathlength to within 50 microns and this took us awhile to find. Later in the night we found it. We then were able to successfully close the pahtlength loop. This milestone represents the first end-to-end closed loop verification of the LBTI system. Elwood and Denis worked on refining the loop parameters. In open loop the system showed >600 nm RMS pathlength variation. In closed loop we were able to consistently reduce it to 300 nm.


Plot of the pathlength RMS over a couple minutes. The "bump" in the center is when the loop is open. Times 0-70s and 100-130 s are when the loop is closed.

2 comments

December 2013 Run: Day 1

The LBTI group is back at the summit for our December observing block.

We've had several gremlins pop up at the beginning of the run:
  • Last weekend Manny uncovered a power supply failure on the right side. This is similar to the one seen on the left side in October 2012. It suggests we need to redesign the DC power in the racks. Paul and Mitch replaced the power supply on Tuesday.
  • In the afternoon Vanessa solved a failure of the WFS on the right side by replacing an A/D board. This appeared to be caused by the power supply failure.
  • The LBTI webserver computer had a corrupt disk. Elwood repaired and checked for any damaged files.
  • The AO software had a few startup glitches, but now appears to be running fine.
  • The espresso machine has failed. However, Vanessa and Andy have figured out how to use a percolator, so we may be okay.
It was windy at sunset. Andy bet Phil that we would be open for the first night. As it turned out we did open for 15 minutes at the beginning before winds forced us to close. Phil maintains that this doesn't count. Andy refuses to pay up.

In the meantime, we are continuing work with the artificial star source to test the phase sensor in poor weather conditions.


1 comment

October 2013 Run Day 4: Phasing Tests on Several Stars

On Friday October 18 we had a half night scheduled for LMIRCam science, followed by a half night for phase sensing engineering. The LMIRCam observations went smoothly and had good conditions.

For phase engineering, we saw significantly more phase variation with our revised loop algorithm (see last post). The phase variations appeared to be nearly +/- 180 degrees per frame at 2 microns. This suggests a vibration problem. We tried shutting off
  • our own compressor, turning off
  • the mirror ventilation
  • the facility chiller
  • turning down the AO gain.
None of these seemed to affect the phase variations. So, we took as much data we could. We acquired:
  • 10 micron nulling data at 30 Hz (see plot from Denis Defrere below)
  • 2 micron phase data at 500 Hz (data still being analyzed)
  • 4 micron dispersed fringe data at 3 Hz (shows large pathlength motions)
As the plot shows there seems to be some vibration at 10 Hz, but 200 nm is about 1/10 of a wave at 2 microns, while the 2 micron data (the error signal for our control loop) seemed to be showing up to 500-1000 nm variations. For the moment, it's a mystery what's going on until we complete the data analysis.


Saturday through Monday observations are LMIRCam science, so we have a couple days to determine our next step for phase sensing. We will resume tests on Tuesday night.


2 comments

October 2013 Run Day 3.5: Closed Phase Loop Off Sky

After puzzling over last night's (UT Oct. 18) data we realized the phase signal was not being reliably calculated. Elwood scrubbed the code, and the afternoon of Oct. 18 we were able to show that we could control the pathlength of the interfereometer via the 2 um fringes.

This is a video of Elwood demonstrating the loop. We are looking at the FFT image of the 2um detector. The plot (barely visible) is showing the angle of the spots on the screen. As we are watching it, he changes it from 0 to 90 degrees.


The artificial source we were using showed 100-300 nm RMS pathlength variations. This suggests something is creating vibrating in our instrument. While this will not hold up further engineering, it is certainly something we need to fix before regular interferometric observations.

We also notice a drift of ~100-200 um in pathlength difference as the telescope slews from horizon to zenith.



5 comments

October 2013 Run: Days 2 and 3

The nights starting on October 16 and 17 were split evenly: first half science projects, second half phase sensing tests.

On October 16 we began by aligning and testing out the new AGPM coronagraph in LMIRCam. Denis Defrere obtained good quality data using the device after careful alignment by Denis and Andy. In parallel we used PISCES to take shorter wavelength observations with the right side telescope.

Caption: The AGPM being aligned. The central hole in the star shows the suppression of the stellar image.



We have a new graduate student who joined the group this fall: Amali Vaz. Although her primary area is helping out with the phase sensing and correction subsystem (Phasecam), she is learning the ropes of running the AO system this run.


Caption: Vanessa shows Amali how to run the left AO system. The irght AO system is running smoothly on the computers to Vanessa's right.



AO operation was so smooth, that Amali entertained herself making TINY cranes.


On the second half of October 16 we had to solve the puzzle of why we saw only poor quality 10 um fringes the first night. After extensive searching we realized that an internal alignment in NIC over the summer changed the setpoint of the nulling interferometer by ~1.2 mm. Once this was achieved we saw good quality interference (the interference last night must have been due to a ghost image). We spent the rest of the night fine-tuning the tip-tilt feedback portion of Phasecam and taking tests open loop data.



The fringe quality and stability started out good and degraded toward the end of the night.


The afternoon of October 17 we used our internal artificial source to check loop operation. This is a new addition to NIC (well newly aligned anyway) that will make nulling and phasecam tests offline much more straightforward.

The evening of October 17 we started out with some LEECH observations. At midnight we again transitioned to Phasecam testing. After our off-sky testing we could reliably align and setup Phasecam. Tip-Til stability was confirmed. We then took phase sensing measurements. While some improvement may have been seen, we realized further refinement with the off-sky source is the best path forward. We handed the telescope back to the LEECH program for the last couple hours of the night.

0 comments

DX AO - new and improved

Just now getting around to writing a post about our AO recalibration from last week, so these posts will be a little out of order.

With the new adaptive secondary mirror (ASM) installed on the right side (DX side), we're nearly ready for binocular (dual-aperture) operation again (see the LBTO ASM blog for all the ASM re-commissioning details). Our complete adaptive optics system has two main components: the ASM and the wavefront sensor (WFS). The ASM can flex to correct for image distortions injected by the atmosphere, but only if it is provided with a measurement of those distortions. These measurements come from the wavefront sensor, so-called because the technical term for these types of image distortions is "wavefront distortions," or distortions in the light waves that reach the telescope. The WFS is essentially a camera which uses a clever way of imaging a star to sense the distortions to that star's light.

Over the last month, Guido, Phil, and I have been preparing the wavefront sensor hardware and software for the calibration process, ironing out the kinks that were preventing us from reaching the best performance in the past. The images on the wavefront sensor must have a precise magnification and geometry in order to provide accurate wavefront measurements, so we've been playing with some of the hardware within the WFS unit to improve our images.

The WFS and ASM must be calibrated together in order to form a working system, which means that because we have a new mirror and a re-aligned WFS, we need a new calibration.

Jared, Guido, and I arrived at LBT on Thursday to finish up the process. Alas, Jared had to leave his nunchuck collection at home.









We were greeted by howling winds at the summit - 90mph gusts the night before had shattered one of the windows in the common area (luckily only the outer pane of a two pane window).






























All the wind and fog iced over the dome, unfortunately for the observers. Thankfully, it had all melted off by the end of the next day.







Inside, the dome was much calmer, and we got to work aligning the calibration unit.














During the calibration process, we put known shapes (called "modes") on the mirror and measure the resulting image distortions with the wavefront sensor. The more modes you can sense and correct, the better your final image will be. Thankfully the Arcetri team had laid the framework for us, providing a validated set of 500 modes for us to use for calibration.

Over the course of the next several days, we took closed-dome (off-sky) calibrations with up to 500 modes - the highest number of modes we've ever been able to calibrate on either side!


While we don't yet know if the 500 mode calibration can be stable on a real star (the seeing was too poor to use it last night), we our new 400 mode calibration performed well on-sky. For the first time, both the right and left sides have 400 modes of correction available, which should make for some excellent high-resolution images this fall.


quote of the day(s): "That's a great modal plot!" - Alfio, approving of the performance of our 500 mode calibration. We breathed a sigh of relief after that!
[He was referring to the plot below, created closed-dome using simulated turbulence. The x axis is the mode number (low spatial frequencies on the left, high on the right) and the y axis is the wavefront error in each mode. The red line is the raw wavefront, and the white line is after AO correction]


1 comment

October Run: Day 1

The evening of October 15 was our first night on the sky for the Fall 2013 semester. Manny, Mitch and Phil spent the afternoon installing the SPC and aligning the right arm of the beamcombiner. This was our first night on the sky with the new shell on the right telescope. We had also carried out a complete realignment and addition of mechanisms for the nulling channel. The night had clear conditions and variable seeing (1-3 arcsec).

Highlights of the night:
  • Both AO systems were operating with 400 modes correction by approximately 1.5 hours after sunset.
  • We were able to scan for interferance and find the new setpoint by 10 pm. (The SPC had been removed).
  • NOMIC-Nulling and LMIRCam images were coaligned for the first time. (allowed simultaneous interference measurements).
  • Dual Telescope Offset Pointing was debugged and demonstrated. This allows automated nodding.
  • The new PZT software was tested out and used.
  • We tested out the UBC alignment motors.
  • We tested out the AGPM coronagraph and refined the alignment procedure.

We did have a few problems:
  • NOMIC communication was flaky throughout the night.
  • A curved slit in LMIRCam was installed incorrectly and needs to be fixed for that science program.
  • We could not detect interference with Phasecam despite multiple scans.
  • Nulling at 10 microns was almost nondetectable.
We think the last two were due to poor atmospheric seeing and will try under what are supposed to be better conditions tonight.

In all, it was a good first night.

0 comments

Presentation at OHP conference (France)

Denis Defrère was at the conference "Improving the performance of current optical interferometers & future design" taking place at the "Observartoire de Haute Provence" (OHP, France). Here are the slides of the presentation he gave on the LBTI.

Download file "Defrere_OHP2013.pdf"



0 comments

NASA Site Visit, Sept 5

Phil hosted (no pun intended) a NASA site visit last Wednesday and Thursday. Wednesday was review at Steward while Thursday was a visit to the LBTO. Hashima Hasan and Gary Blackwood had yet to see the telescope and instrument so it was quite a treat for us to show them around... Here are a few pictures. The rest can be found at http://lbti.as.arizona.edu/gallery.



l-r: Ben Parvin, Hashima Hasan, Phil Hinz, Rafael Millan-Gabet, Gary Blackwood, Mario Perez
Hashima Hasan and Ben Parvin




Phil Hinz, Hashima Hasan, Mario Perez, Gary Blackwood on Level 9 platform.
This is the image that Phil really shows his aura...

A good trip in all....

6 comments

July LBTI Observing. Now With More Clouds!

We're on night #3 of an LBTO observing run, but the heavy monsoon activity has kept us closed the whole time. The group this time around is Denis and myself, along with visiting Science Team member, Bill Danchi, and summer student Jon Kyl.

Elwood and Paul visited us and were able to troubleshoot some new software for PZT control. Denis has been testing NOMIC, but not much else is able to be done in these conditions.

So, we took a walk the past couple afternoons. It is definitely more pleasant on Mt. Graham this time of year than in Tucson.



As pleasant as the walks are, when we get back to the observatory, we see this:


So we've been occupying ourselves with movies and basketball games in the evening.




1 comment